11 11 TEX L^{A}TEX

Wearable-HOWTO.

Lionel, trollhunter Bouchpan-Lerust-Juéry <trollhunter@linuxfr.org> v0.0.5, December 1999

Nomadism is one of the major trends of our society, now most of the people can work with a laptop computer, but few did the next step: to live, or if you prefer to wear one computer that is a wearable.

Contents

1 Foreword 2

2 What is a Wearable anyway ? 2

3 Copyright, Disclaimer and Trademarks 3

4 Advocacy 3 4.1 Why ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 4.2 The Wearable concept is not a revolution, it is just an evolution. . . . . . . . . . . . . . . . . 3 4.3 The Wearable may give its user an edge. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.4 The Wearable concept: a stealth menace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.5 Wearable computing: an emerging industry. . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

5 What CPU ? 4 5.1 PDA based Wearables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 5.2 PC/104 and laptop based Wearables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

5.2.1 What is PC/104 ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5.2.2 What about PC/104-plus ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

5.3 Misc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

6 Power supply 5

7 OS 5

8 The Sulawesi project. 5

9 Input 6

9.1 Variations around the keyboard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 9.1.1 Mini keyboards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 9.1.2 One hand keyboard. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

10 Audio Output 6

11 Visual Output 7

11.1 Head Down displays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 11.1.1 Wristwatch type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 11.1.2 PDA type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

11.2 Head UP displays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 11.2.1 Obstrusives displays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 11.2.2 Nonobstrusives displays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

12 Comms 7

13 How can I have my Wearable ? 8 13.1 Commercial Solutions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 13.2 Do it yourself. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

14 PalmPilot and its clone ( IBM, HandSpring, TRG ): a new breed of wearables. 8

15 How to carry my wearable ? 9 16 Applications with Wearables. 9

16.1 In the army now. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

16.2 At the hospital . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

16.3 With the firefighters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

16.4 Wearable for the disabled. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

17 A borg’s life. 11

18 Sources of Information 11

18.1 Non commercial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

18.2 Commercial. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

19 To do List. 12

20 Revision History 12

21 Thanks and Credits 12

1 Foreword

I am not a guru in wearable computing, simply after some years using HP95LX and HP200LX palmtop computers more like wearables than like a PDA ( thanks to DOS 5, the HP200LX even allowed me to do a semester CS project using turbo pascal while commuting in the subway ), and living with a palm III and a laptop running Linux, to me the next step is having a wearable running Linux, thus I have spent more than a year reading various informations on wearable and taking a little part in the wear-hard mailing list. As more an more people will be interested in the wearable concept I have decided to start this HOWTO in order to help them ( At the time of this writing I am building my own wearable using my laptop ( Toshiba 430CDT ) with a twiddler keyboard, emacspeak, and camcorders batteries. )

Some people may, at least, be surprised to find the "In the army now" section so I would like to explain my position: I am trying to write this HOWTO both as a journalist and a scientist would do: this implies as a journalist to look for information from every source available, and as a scientist to be neutral: and as the US Army is willing to purchase more than 30.000 units one can hardly overlook this user group. This document is intended to be the starting point of a long term work. If you have any information ( that can be checked ) that you think is of interest, do not hesitate contact me.

2 What is a Wearable anyway ?

Well, as it name implies a wearable is a computer you are supposed to wear, actually very few wearables fill this definition. On a more practical point of view one can define a wearable as a computer you can use on the run relying only on its power supply ( see the power supply section for more information ). So with such a definition PDAs, palmcomputers, customized laptops are wearables too.

I am not a specialist in wearables, simply I have a very strong interest in nomadic computing and I studying a lot its implementations both from a technical point of view and from the way nomadic computing is going to change our every day life and the way we interacts as we are getting more and more connected.

This text is included in the LINUX DOCUMENTATION PROJECT http://linuxdoc.org/LDP <http://

linuxdoc.org/LDP> .

The latest version of this document is available in HTML format at http://infonomade.linuxfr.org/Wearable-

HOWTO.html <http://infonomade.linuxfr.org/wearable/Wearable-HOWTO.html>, you can retrieve the

sgml format at http://infonomade.linuxfr.org/Wearable-HOWTO.sgml <http://infonoamde.linuxfr.

org/wearable/Wearable-HOWTO.html>

I will maintain the French version too. If you wish to mirror it or to translate it, please contact me.

Lionel, Trollhunter Bouchpan-Lerust-Juéry <trollhunter@linuxfr.org> or at <boucli27@altavista.net>

3 Copyright, Disclaimer and Trademarks

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999 by Lionel, trollhunter Bouchpan-Lerust-Juéry. This document may be distributed under the terms set forth in the LDP license at http://linuxdoc.org/LDP/COPYRIGHT.html <http:

//www.linuxdoc.org/LDP/COPYRIGHT.html> .

This is free documentation. It is distributed in the hope that it will be useful, but without any warranty. The information in this document is correct to the best of my knowledge, but there’s a always a chance I’ve made some mistakes, so don’t follow everything too blindly, especially if it seems wrong. Nothing here should have a detrimental effect on your computer, but just in case I take no responsibility for any damages incurred from the use of the information contained herein.

In this document you will encounter some commercial products and brands. Theses products are cited for information purpose, it is not an endorsement from the author. The trademarks belong to their respective owners.

4 Advocacy

Some people may want to know why you want to build a wearable or you may have to sell your idea to the people who are responsible for the purchase. Here are some ideas you may use.

4.1 Why ?

Because, you want to experiment by yourself, because you think that Human Computer Interaction is not a matter of designing the n-th windowmanager, because you feel that enhanced reality is the cutting edge of your evolution.

4.2 The Wearable concept is not a revolution, it is just an evolution.

Some people are afraid of revolutionary or so called products, thus the good news is that the wearable may look like a revolutionary concept but it is actually just an evolution of the computing hardware. Let me explain that : At first there was the Mainframe, then came the desktop computer enabling people to work in their office, latter the PC enabled these people to work at home too, as time passed the portable PC enabled people to work in a Hotel room, or everywhere they could find a power plug ( yes an Osborne or an IBM’s convertible were definitively not laptop computers ), at the same time some pocket computers appeared on the market : Sharp PC1500, Canon X07, Casio PB100 then the first one weighted less and less, the second one disappeared but the LCD screen was, with other things their legacy allowing the laptop to emerge, as the laptop went mainstream, its size went smaller allowing people to work in the train, at the library ..., then palmtop PC’s such as the HP95LX and PDA’s appeared, ( at this time the most successful is the PalmPilot family and its clones ) allowing people to work on the move, so the wearable is just the next step in this move towards miniaturization. ( If you want to learn more about Laptops and Linux you should read the Laptop-HOWTO, the latest version can be found at Werner’s Heuser Homepage

<http://www.snafu.de/~wehe/index_li.html> )

4.3 The Wearable may give its user an edge.

In today’s competitive world it is very important to get an edge over the other company, thus for example in a plane repair company the engineers who are using wearables do not waste time in asking for blueprints but instead have the blueprint and the technical data at will while performing their job, thus they will be able to repair the planes faster. You may choose an example in your job.

4.4 The Wearable concept: a stealth menace.

When they made their first appearance on the market, some products or technologies were, to say the least, less than perfect. This is a stealth menace, if you read Clayton Christensen’s book "The innovator’s dilemma: when new technologies cause great firms to fail" or Andy’s Grove "Only the paranoid survives" you will see that some corporations that relies on a product and that have a king of the hill may fall down because they overlooked a new product/technology that was clumsy at its beginning and was at first in a market niche, then the contender took over the market.

4.5 Wearable computing: an emerging industry.

The people who decide to fund your project are not always very found about technical details, thus you will have to use other arguments, otherwise your pet project will be sent to /dev/null. If one reads again chapter 2 of AP Sloan’s book: "My years with General Motors" it is obvious that the Wearable industry is going into the same changes as did the automobile industry in the US at the turn of the century : it turned from a Hobby with some small manufacturers to a mass market Industry. It took nearly 30 years to the automobile industry to change, but the wearable industry should have done this changes in less than 5 years, so if they don’t invest on this product others will do.

5 What CPU ?

5.1 PDA based Wearables

Usually they use non Intel CPU ( except the HP95/100/200LX ) See the PalmPilot : a new breed of wearables section.

5.2 PC/104 and laptop based Wearables.

Theses devices are using usually an Intel or Intel compatible CPU.

5.2.1 What is PC/104 ?

PC/104 is an industrial standard, for PC based systems that can be stacked together in order to create an embedded system. The footprint of theses boards is 4"x4" so as you can see it is a good base in order to start a wearable project.

5.2.2 What about PC/104-plus ?

It is a PCI addition to the PC/104 standard. you can read the PC-104 FAQ <http://www.controlled.

com/pc104faq>

5.3 Misc

Last year there were some alpha based Multia sold for around $100, but it seemed AFAIK that nobody ever managed to build a wearable around one of theses devices.

6 Power supply

There is plenty of choice as long as you use batteries. But be careful __NEVER__use liquid acid filled batteries it is a Health Hazard for this application.

7 OS

For the people whose Wearable is based on a PDA, there is little choice : most of the time they have to do with the PDA’s manufacturer OS, AFAIK Linux on Psion series 5 is not mainstream. a commercial solution called WindStone primary targeted to palmpilots with a project to port it to ARM based computers it is

made by OSK <http://www.oski.co.kr/>, so wait and see

For those using a PC derivative, the choice seems to be between Microsoft products or Linux, Well, Microsoft products are not exactly the best choice and DEBIAN/GNU Linux is the winner ( I don’t think that for general purposes one has to use a posix .4 compliant kernel ).

There is another contender QNX <http://www.qnx.com>, but is more industrial oriented, so for hobby or research it is not the best choice.

8 The Sulawesi project.

This project deals as well with input modalities and with output modalities thus its section is just before the Input/Output part of this document.

It is meant to accept input from different modalities and to do its input to a number of different modalities, this is written in Java; The good news is that it is highly portable, the bad news is that today a JVM is a memory and resource hog ( but, to me, Java and the related technologies such as Jini, just like the wearable, is a stealth menace ). One of the nice features of this project is the fact that, because of a clever design, it can deal with with a great input and output modalities that exists today, and when new modalities one hadn’t even think about will be available, it should be able to deal with them in a very short time and with little effort.

Today, can accept input from a GPS too. You can find more information at : Sulawesi Home Page <http:

//vasewww.essex.ac.uk/projects/>

9 Input

9.1 Variations around the keyboard

9.1.1 Mini keyboards

Usually theses devices are on the wearer’s forearm or wrist.

The WristPC <http://www.pfuca.com/products/hhkb/hhcindex.html> Keyboard and the Arm mount

micro keyboard <http://www.ivpgi.com/1accessory.htm> from the Phoenix Group Inc. are typical from this concept.

9.1.2 One hand keyboard.

IMHO it is the best concept because it doesn’t require you to look at the keyboard while typing is the less intrusive when you have to deal with other people, and in order to be accepted by other a Wearable wearer’s must have the lowest visual signature.

The archetype of this concept is the Twiddler <http://www.handkey.com> and all of its clones.

10 Audio Output

When your system doesn’t output too much data, or in order to signal some special events an audio output can do the job. Most laptops come with a Sound Blaster compatible sound card and a lot of PC-104 cards are sold with such a sound card on board or you should buy a sound module. The next step is to redirect your text output to a text_to_speak program.

As a lot of input is done using emacs, it seems to be a good idea to use Emacsspeak <http://www.cs.

cornell.edu/home/raman/emacspeak/emacspeak.html>

11 Visual Output

11.1 Head Down displays.

11.1.1 Wristwatch type

Some manufacturers ( eg Seiko, Citizen, Casio ) manufactured Wristwatches that could be linked to a computer, ( I remember that in 1982 Seiko showed a prototype of a wirstwatch TV ( monochrome LCD )).

It is also possible to use a 4 lines LCD display connected to the serial port of your laptop ( numerous paper or online electronics magazine offer schematics ).

11.1.2 PDA type

PDA’s seem to be a cheap and efficient way of sending and displaying data for a wearable, actually any PDA with a VT100 emulation program and a serial link can be used effectively as a terminal ( I have successfully used my HP200LX running

kermit <http://www.columbia.edu/kermit>

as a terminal for my desktop, giving me an emergency access to it if the display failed )

11.2 Head UP displays

The distinction between obstrusives an non obstrusives ones is not from the wearer point of view but from the other people, that is is the display forbid to see the wearer’s eyes

11.2.1 Obstrusives displays

Hacks with LEDs Some people managed to use blinking LEDs in order to retrieve information from their device. This is one of the simplest display one can imagine : a LED or a row of LED blinking.

Sony GLASSTRON Sony sells a device called the Glasstron

The VRD ( Virtual Retinal Display ) With this device a LASER LED draws the image on the wearer’s retina, the US Navy tested it in the summer 1999, at Hawaii ( see in the army now ).

11.2.2 Nonobstrusives displays

The Microptical <http://www.microopticalcorp.com>corporation manufactures two displays the Integrated Eyeglass display and the ClipOn display.

12 Comms

AFAIK Most of the time it is with a radio link, should it be with a radio transceiver or via a cell phone.

13 How can I have my Wearable ?

13.1 Commercial Solutions.

Some companies sells Wearables ready to use out of the Box, hereafter is a list of them.

   • Genesys Technology <http://www.genesistech2000.com/product.html> manufactures a Wearable

that uses a HMD display.

• handsfreemobile <http://www.handsfreemobile.com> sells the Mid Riff Brain, it is mostly an LCD touchpad based solution and the unit is in a pouch fixed on the waistband.

   • Teltronics <http://www.teltronics.com/is/hardware.html> sells the Mentis. It is a modular wear-

able.

• Via <http://www.flexipc.com/> sells the VIA II PC a Wearable that is worn on the waistband.

• Xybernaut <http://www.xybernaut.com/> sells the Mobile Assistant IV ( MA IV )

13.2 Do it yourself.

As wearable computing is in its very early beginning you should make various experiments and share them with your peers. Doing this with an open and versatile OS such as Linux should give you a lot of fun ( after all, desktop computer is just a plain vanilla computer), because you play with both the OS and the hardware. If you are unsure of the technical decisions you are going to make, you can start with a proven design such as the

Lizzi <http://wearables.www.media.mit.edu/projects/wearables> design and then change some com-

ponents. Thus you will not have to reinvent the wheel and you will avoid some pitfalls.

14 PalmPilot and its clone ( IBM, HandSpring, TRG ): a new breed of wearables.

The PalmPilot family and its clone is a hit among PDAs, but less known is the fact that it is more than a PDA it is one of the first off the shelves wearable. You can connect it to your cell phone, When on the move you may need an input device more convenient than Graffiti: a keyboard for example, one device exists that allow you to plug a PS/2 keyboard in your palm: Happy Hacking cradle Specially for the Palm.

PFU <http://www.pfuca.com/products/hhkb/hhcindex.html> Happy Hacking sells a PS/2 adapter for

the Palm, thus one should be able to hook any PS/2 keyboard on this device. It is battery ( 2 AA ) powered.

It seems that someone has managed to get a twiddler keyboard working with the PalmPilot. If you want to do land navigation it exists an electronic compass you can plug into your palm it is called Palm Navigator and it is manufactured by Precision Navigation <http://www.precisionnav.com> They even sell a weather station.

If you have more money to spend or want more precision you can purchase a GPS receiver called Earthmate; it is manufactured by Precision Navigation <http://www.delorme.com> Some of you may want a software only solution: this can be done. If you have some highly specialized needs like commuting without having the subway map and want a readily made solution you can use a program like Route Expert, I use it on a daily basis, it works great with Paris’s subway and I will test it with Berlin’s subway at the end of the year ( of course it exists other programs but this one is the one I am using ). I guess everyone can find the North using the sun and a wirtswatch, a PalmPilot can do it with more precision using readymade programs such as T.J’s Sun-Compass ( once again there exists other programs but this one is the one I am using ).

The most important with the PalmPilot for the Linux community is the fact that there exists two ports of Linux for it ( DragonBall is a member of the 68000 family after all ). One is GPLed

You can find more information at : uClinux <http://www.uclinux.org>

the second one is commercial, it is based in Korea. WindStone <http://www.oski.co.kr> it is based on uClinux and it is coming with its own GUI and widgets.

Thus with the next generation of Palms, the CPU speed will be twice today’s speed, and with Linux on board we will be able to run a lot of software that are hardware independent and have been written for today’s x86 Linux based wearables.

15 How to carry my wearable ?

Depending on your wearable’s family ( eg Palm, palmtop computer, laptop based, PC/104 ) there seem to be 3 convenient ways of carrying it around. You can carry it on tour wrist, or in a backpack or you can hang it on your waistband.

16 Applications with Wearables.

As a rule of thumb the first people who are going to use extensively the wearables are the people whose work is a matter of life and death.

16.1 In the army now.

It is very important in an army to have an edge other the enemy, that is what we have always seen in history, and moreover the army is known to spend vast amount of money in order to get this edge.

Some years ago the infantryman was ordered to carry different pieces of ordnance that came from different sources and were not part of a whole thing, and to try to do his best with theses ( looks like Frankenstein to me ). With theses new projects the infantryman is the centerpiece of a coherent, complementary weapons and equipment system.

• French Forces. In France there is a project called FELIN ( Fantassin à Equipement et Liaison Integrées ( Foot soldier with integrated equipment and links ) ) in order to enhance the efficiency of the soldier a set of devices had been developed with very careful limitations : historically the foot soldier has been a beast of burden, thus the FELIN project wishes to limit its load at 25kg. The soldier has got a radio link, a computer and a camera on his weapon ( allowing him to open fire from cover ). The display is done with a monocular HMD.

• UK Forces. The FIST ( Future Integrated Soldier Technology ) project: Early tests were conducted on Salisbury Plain to determine if the soldier could access information either ’Head Up’ ( in a helmet display ), ’Head Down’ ( on a wrist mounted display ), via a palmtop computer or if necessary a map lessons learned ( in L’armement issue 67 September 1999 )

– There is no scope for enhancing soldier performance through the provision of tactical information unless the soldier can access information ’on the move’

– The preference, both by day and night was the helmet display.

The FIST Digitization Trials will be conducted on Salisbury Plain in June and November 2000. The June trial addresses information flow, the November trial will address the benefits of enhanced information to determine if there is a consequent increase in the ’Tempo of Operation’.

• US Forces

– The land warrior program It is a system made of 5 subsystems cooperating together.

∗ A Pentium Computer/Radio Subsystem (C/RS) made by Motorola featuring an integrated GPS receiver

∗ The Protective Clothing and Individual Equipment Subsystem (PCIE) made by Gentex ∗ The Weapon Subsystem (WS) made by Raytheon

∗ The Software Subsystem made by Raytheon

∗ The Integrated Helmet Assembly Subsystem (IHAS) made by Honeywell The location of each squad member will be available through the IHAS, as well as digitized maps and tactical information. Every soldier will have an integrated GPS as well as video recording capabilities. Of course the soldier will have night vision capability as well as deported sighting. Tests should take place during year 2000. More info on the Land Warrior program at http://www.sbccom.army.mil/programs/lw/index.htm <http://www.sbccom.army.mil/

programs/lw/index.htm>

( There is even a FAQ and pictures ). AFAIK this is going to be the biggest test for the wearable concept because the Army is planning to purchase 34000 units plus spares.

– The navy. During the summer 1999 news poured that the US Navy tested the Virtual Retinal Display at Hawaii. The explanation is that there doesn’t seem to be enough space in the fighting ships because of the numerous CRT, so it is conceivable for the crew to wear VRDs Needless to say, that some very low profile and highly specialized units should have adopted the wearable concept but as far as we cannot have evidence we can only speculate on that.

16.2 At the hospital

The Microvison <http://www.mvis.com>’s VRD has been delivered <http://www.prnewswire.com/cnoc/

exec/menu?127059> to the Wallace Kettering Neuroscience Institute, Dayton Ohio, for neurosurgery applications. This is not a full featured wearable solution but it is a big step toward the adoption of wearables in health care activity.

16.3 With the firefighters

Firefighters are using thermal cameras in order to see through the smoke, in the next years theses devices should be miniaturized and with a wearable firefighters should have both hands free and will have real time access to data such as the map of the building and numerous helpful informations.

16.4 Wearable for the disabled.

One can easily think that a wearable computer can be very helpfull for the disabled people, for example for blind people a wearable with a GPS receiver, some maps of the surroundings and even more, interacting with its surroundings by means of active beacons can be a very good substitute for a dog, the elements exists for a long time.

This should be easily done : the linux comunity has good text to speech applications, voice recognition is quite ready for this application , the IrDA or Modem radio drivers are OK, the same applies to the GPS programs, we have to develop an electronic blind walking stick in order to detect the obstacles at a much reater range and last, but not the least to convince the autorities.

17 A borg’s life.

This section deals with Human-Human interaction. As the wearable is quite new for the layman one has to expect some strange reactions from the other people. When I go out and read my PalmPilot’s screen while walking or in the subway people have one of the following reactions :

• No reaction: because they are in their own world or because they don’t care.

• Curiosity: they come to me and ask me questions like "what it is", "what are you doing with it" and "how much does it cost ?"

• Disdain : they may think I want to draw attention.

• Hostility/Fear : "this guy is nut: avoid him".

Steve Mann Gave an interview in the New Scientist magazine and as he was one of the first to test the waters, so he has a long experience in this field. The second problem is that the wearable’s hardware is quite expensive and some predators will spot a new device that can be easily stolen and sold at a good price, or just in order to have one of them. As a conclusion you have to be very cautious.

18 Sources of Information

18.1 Non commercial

A good starting point to gather information is to subscribe to the wear-hard@haven.org mailing list, you have to send a message with the word subscribe in the subject to wear-hard-request@haven.org.

The archive of this mailing list is mirrored at Wearable Central <http://wearables.blu.org> This site was founded to be an archive center for the mailing list and the newsgroup. It is maintened by R. Paul McCarty.

18.2 Commercial.

You can find information on the web sites of the company cited in the various sections of this document and as stated earlier in this document : The companies quoted here are just for your information only, I do not endorse any of their product, this just in order to help you.

19 To do List.

This section is about things that have to be done for the next releases if you wish to contribute it is a good place to look at.

• To review more in depth the sulawesi project.

• Organic Leds are very interesting, I should cover theses too.

• To expand the input/visual output sections.

• To expand and reorder the advocacy section.

• To have an "in the car" section.

• The concept of an UW wearable seems viable lets think about it.

• List and describe all projects.

• List and describe all products.

• To get some sleep.

20 Revision History

• v 0.0.5 December 1999, first release.

• v 0.0.1 November 1999, first draft.

21 Thanks and Credits

I would like to thank among other people:

• Werner Heuser <wehe@snafu.de> <mailto:wehe@snafu.de> for the work he has done maintaining his

  Laptop-HOWTO <http://www.snafu.de/~wehe/index_li.html>

• The people of the Wearable mailing list which proved to be very informative and open minded.

• My friend Manu for his never ending patience and kindness.

• The people of the Wearable community whose discussion and web sites strengthen my confidence in the wearable concept.